how do i solve the multi-step equation x-2(3x-2)=-6?

Answer by tjnw79(57)

x-2(3x-2)=-6
first times everything by -2 in the parenthsis
x-6x+4=-6
now subtract the 4 on both sides
x-6x=-10
add you like terms
-5x=-10
now divide by -5 on both sides to get
x=-10/5 or x=2